OpenStudy (anonymous):

evaluate factorial expression 14!/2!12!

OpenStudy (dls):

\[\Huge 2!=2\]

OpenStudy (dls):

\[\Huge \frac{14!}{2! 2!}=\frac{14!}{4}\]

OpenStudy (anonymous):

Do you know what ! means?

OpenStudy (anonymous):

ans is 13*14/2

OpenStudy (anonymous):

you can cancel 14! and 12!..so what is left of 14! is 13*14....and 2!=2.whenever you see factorials try cancelling them out.otherwise,it can get messy

OpenStudy (psymon):

\[\frac{ 14*13*12*11*10*9*8*7*6*5*4*3*2*1 }{ (2*1)(12*11*10*9*8*7*6*5*4*3*2*1) }\] Factorial means the multiplication of that number by every number counting down all the way to 1. If you wrote it out (not always practical) like this you can see that the whole 12! cancels leaving just (14*13)/(2). Everyone is saying the correct stuff, this is just the visual of what is there.

OpenStudy (anonymous):

\[\frac{ 14*13*12*11*10*9*8*7*6*5*4*3*2*1 }{ 2*1*2*1 }\]